4. Find the probability of randomly selecting 4 good quarts of milk in succession from a cooler containing 25 quarts of which 5 have spoiled, by using:

(a) Theorem 3 of the product rule
(b) the concept of probability of events
1
Expert's answer
2020-11-04T16:40:31-0500

There are 25 quarts of which 5 are spoiled. Hence there are 20 good quarts.


a) product rule:

P(4 good quarts) = P(1st is good AND 2nd is good AND 3rd is good AND 4th is good) =

= P(1st is good) * P(2nd is good) * P(3rd is good) * P(4th is good) =

= "\frac{20}{25}\cdot\frac{19}{24}\cdot\frac{18}{23}\cdot\frac{17}{22}=0.38"


b) the concept of probability of events:

Number of selecting 4 good quarts is combination:

"C(20,4)=\frac{20!}{4!16!}=4845"

Number all possible outcomes selecting 4 quarts:

"C(25,4) =\frac{25!}{4!21!}=12650"

P(4 good quarts) = "\frac{C(20,4)}{C(25,4)}=\frac{4845}{12650}= 0.38"